I do not get the following questions:
y-intercept of -2.1,x-intercept of 3.5


through (1.2, 5.1), x-intercept of 3.7

Answer by Alan3354(69443)

y-intercept of -2.1,x-intercept of 3.5
At the y-intercept, x is zero, and vv.
So 2 points are given, (0,-2.1) and (3.5,0)
The slope is diffy/diffx = -2.1/3.5
m = -0.6
Since the y-intercept is known, y = mx + b
y = -0.6x + 3.5
-------------------
through (1.2,5.1), x-intercept of 3.7
The x-intercept is (3.7,0)
m = (0-5.1)/(3.7-1.2) = -5.1/2.5
m = -51/25 = -2.04
Then use
y-y1 = m*(x-x1) where (x1,y1) is either point
y = -2.04(x - 3.7)
y = -2.04x + 7.548